OpenStudy (anonymous):

PLEASE ASAP Find the vertex, focus, directrix, and focal width of the parabola. -(1/12)x^2 = y Vertex: (0, 0); Focus: (0, -3); Directrix: y = 3; Focal width: 48 Vertex: (0, 0); Focus: (-6, 0); Directrix: x = 3; Focal width: 48 Vertex: (0, 0); Focus: (0, 3); Directrix: y = -3; Focal width: 3 Vertex: (0, 0); Focus: (0, -3); Directrix: y = 3; Focal width: 12

OpenStudy (psymon):

Well, the equation is of the form x^2 = 4py. Notice how I put x^2 by itself, though. That's one tip to start off with : )

OpenStudy (anonymous):

so is it A?

OpenStudy (psymon):

Focal width is equal to 4p. I know the temptation is to multiply that -12 by 4, but it already is 4p. So that 12 is your focal width.

OpenStudy (anonymous):

ohhhhhh okay I understand. What about this one: Find an equation in standard form for the hyperbola with vertices at (0, ±4) and asymptotes at y = ±1 divided by 4.x. y squared over 16 minus x squared over 64 = 1 y squared over 16 minus x squared over 256 = 1 y squared over 256 minus x squared over 16 = 1 y squared over 64 minus x squared over 4 = 1 ?????

OpenStudy (psymon):

Well, since the y-coordinates of the verticies are changing, its a y^2/a^2 form where a = the distance from the center to the vertex. So the distance is 4, meaning a^2 = 16. as for the asymptotes, when the hyperbola is in y^2/a^2 form, the asymptotes are at +/- (ax/b). So if a, which is your numerator, is 4, then what does b have to be in order for it to come out to 1/4?

OpenStudy (anonymous):

16?

OpenStudy (psymon):

There ya go. So if b is 16, b^2 has to be 256.

OpenStudy (anonymous):

oh gosh, ty so much

OpenStudy (psymon):

You should maybe keep track of some of these formulas and equations, lol.

OpenStudy (anonymous):

I have them all with me now, I just didn't understand how to start the problem.

OpenStudy (psymon):

Ah. Well,it just gives you basic info. With hyperbolas, the verticies always give you a. Which coordinate changes between the verticies is what tells you y^2 first or x^2 first. You can derive everything from certain basic things given in the problem :P